OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3.

OpenStudy (anonymous):

for this problem, you can simply plug the numbers into point slop formula, which is y – y1 = m(x – x1) where the point is (x1,y1) and your slope is m so you would have y-(-5)=-3(x-2) or y+5=-3x+6 which can be simplified to y=-3x+1

OpenStudy (anonymous):

thank you!